OpenStudy (vlery):

how would I find the exact value of TAN(-765degrees)?

OpenStudy (vlery):

@mathmath333 you know anything about trig?

OpenStudy (mathmath333):

reduce the angle first

OpenStudy (mathmath333):

add 360

OpenStudy (mathmath333):

two times i think

OpenStudy (amoodarya):

hint: tan(360)=0 tan(720)=0 now write 765 as 720+45

OpenStudy (vlery):

I don't get it our teacher was absent today and never gave us a powerpoint to look over

OpenStudy (vlery):

I'll just ask the teacher tomorrow

OpenStudy (mathmath333):

u need to add 360 here \(\large\tt \begin{align} \color{black}{ tan(-765)\\ =tan(360-765)\\ =tan(-405)\\ =tan(360-405)\\ =tan(-45)\\ =-1\\ }\end{align}\)

OpenStudy (mathmath333):

formual is \(tan(2\pi+\theta)=tan\theta\)